OpenStudy (anonymous):

Find the limit of the function algebraically. limit as x approaches negative four of quantity x squared minus sixteen divided by quantity x plus four

OpenStudy (anonymous):

Please provide an equation

OpenStudy (danjs):

\[\lim_{x \rightarrow -4}\frac{ x^2-16 }{ x+4 }=\lim_{x \rightarrow -4}\frac{ (x-4)(x+4) }{ (x+4) }\]